the ratio between an exterior angle and an interior angle is 2:3 find the number of sides in polygon
Answer by Boreal(15235)

The exterior and interior angles add to 180 degrees
2:3
2x+3x=180
5x=180
x=36 degrees
72:108
The exterior angles are 72 degrees; the interior are 108 degrees.
(n-2)/n*180=108; (n-2)/n=108/180; cross-multiply and 108n=180n-360; -72n=-360;n=5
It is a pentagon.